OpenStudy (anonymous):

simplify the radical expression by rationalizing the denominator 2/square root of 42 a. square root of 84/42 b. square root of 42/21 c. 21 square root of 21 d. sq root of 2

OpenStudy (anonymous):

\[a. \frac{ \sqrt{84}}{ 42 }\] b. \[\frac{ \sqrt{42} }{ 21 }\] c. \[21\sqrt{42}\] d. \[\sqrt{2}\]

TheSmartOne (thesmartone):

Multiply sqrt of 42 on both the numerator and denominator.

TheSmartOne (thesmartone):

\(\sf\Huge \frac{2}{\sqrt{42}}\times \frac{\sqrt{42}}{\sqrt{42}}=?\)

OpenStudy (anonymous):

I'm not sure how to solve it

OpenStudy (triciaal):